Allergan is a major pharmaceutical firm. You work for Allergan’s CFO and are evaluating a major and expensive drug trial. The drug trial would require an investment of $95 million today If the trial succeeds, it will be worth $1 billion in a year. You estimate that there is only a 10% chance that the trial succeeds. If it fails, then in a year you will scrap the project and lose the entire $95 million investment. Whether the venture succeeds or fails is independent of general market conditions. The risk-free rate is 3%, the expected return of the market is 12%, and the standard deviation of the market return is 20%. Ignore taxes. What is the NPV of the project?

$775 million

$870 million

$2.1 million

-$95 million
Answer #1

$2.1 millions

Firstly, calculate the appropriate discount rate for project:

As this venture is independent of market conditions thus its Beta would be Zero.

And, as per CAPM, its discount rate would be :

r=r;+ B* (rm-rp)

where,

r = discount rate

rf = risk free rate

rm = market rate

B = beta

putting the values

r=0.03 +0+ (0.12 -0.03)

r=0,03

thus, discount rate = 3%

Now, Calculate the expected value of project in a year

Expected value of Project in one year = $1 billions * .1 + $*0.9

Expected value of Project in one year = $100 millions

Thus,

NPV of Project would be :

= PV of expected value in one year - Initial Cost

NPV = Exp Expected value of Project in one year (1 + 0.03) - Initial cost

100 NPV = (1 + 0.03) - 95

NPV = 97.1 - 95

NPV = 2.1 million

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.